OpenStudy (sammyalabamy):

Write the expression in standard form. 3/(3-12i) Select one: a. (1/-17)+(4/17)i b. (1/17)-(4/17)i c. (1/-17)-(4/17)i d. (1/17)+(4/17)i

OpenStudy (anonymous):

\[\frac{3}{3-12i}\] multiply top and bottom by the conjugate of the denominator

OpenStudy (anonymous):

the conjugate of \(a+bi\) is \(a-bi\) and this works because \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (sammyalabamy):

Thank you.

OpenStudy (anonymous):

\[\frac{3}{3-12i}\times \frac{3+12i}{3-12i}=\frac{3(3+12i)}{3^2+12^2}\]

OpenStudy (anonymous):

the numerator is what you get when you multiply i would leave it in factored form first, because you are going to cancel the three

OpenStudy (anonymous):

which, now that i look, you could have cancelled before starting \[\frac{3}{3-12i}=\frac{1}{1-4i}\] probably would have been easier to start there

OpenStudy (anonymous):

oh , you are welcome

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) satellite73 \[\frac{3}{3-12i}\times \frac{3+12i}{3-12i}=\frac{3(3+12i)}{3^2+12^2}\] \(\color{blue}{\text{End of Quote}}\) looks like there is a typo \[\frac{3}{3-12i}\times \frac{3+12i}{3\color{Red}{+}12i}\] ??